Your answer is partially correct. Try again. For Goody Company, the budgeted cost for one unit of product is direct materials $10, direct labor $20, and manufacturing overhead 80% of direct labor cost. If 25,000 units are expected to be sold at $65 each, what is the budgeted gross profit? Budgeted cost per unit Gross profit per unit Total budgeted gross profit 475,000
Answer #1

A.

Total cost of per unit = Direct Material + Direct labour + Manufacturing overhead

From the question the values are

Direct labour = $20/ Unit

Direct Material = $10/ Unit

Manufacturing OH = 20 * 0.8 = $16 per unit

Total cost of single unit = 20 + 10 + 16 = $46/ Unit

So the budgeted cost per units will be $46

B.

Gross profit per unit = Selling price per unit – the cost of goods sold per unit

                                       = 65 – 46

                                       = $19

Gross profit per unit is $19

C.

Total budgeted gross profit = (Selling price per unit – the cost of goods sold per unit)* Total sale units

                                                   = (65 – 46)* 25000

                                                   = 19 * 25000

                                                   = $475000

Total budgeted gross profit will be $475,000